OpenStudy (anonymous):

What is the arc length if Θ = 7pi/4 and the radius is 5

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dumbcow):

arc length = angle*radius

OpenStudy (anonymous):

\[\theta = \frac{ a }{ r }\]a = arc length r = radius

OpenStudy (anonymous):

so what would i plug in

OpenStudy (dumbcow):

arc length = angle*radius you want to know arc length right? so multiply the angle with the radius...

OpenStudy (anonymous):

whats the angle though? is it \[\frac{ 7\pi }{ 4 }?\]

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yes theta typically refers to the angle 7pi/4 is angle in radians

OpenStudy (anonymous):

\[a = \theta*r \implies \frac{ 7 \pi }{ 4 }(5)\]

OpenStudy (anonymous):

so 27.488?

OpenStudy (anonymous):

or is it \[\frac{ 35\pi }{ 4 }\]?

OpenStudy (dumbcow):

same thing...one is exact the other is a decimal approximation

OpenStudy (anonymous):

so its \[\frac{ 35\pi }{ 4 }??\]

OpenStudy (dumbcow):

depends on what the instructions are regarding the format of answer

OpenStudy (anonymous):

okay. i have another question

OpenStudy (anonymous):

In the supermarket, there is a scale that measures the weight of fruit and vegetables. The needle is 24 cm long and rotated 96° when Charlie placed four tomatoes on it. How far did the needle travel as it rotated?

OpenStudy (anonymous):

1152π 2304π \[\frac{ 5\pi }{ 64 }\] \[\frac{ 64\pi }{ 5 }\]

OpenStudy (dumbcow):

ok they want arc length again note: angle MUST be in radians to find arc length

OpenStudy (anonymous):

so how do i make it into radians

OpenStudy (dumbcow):

pi = 180 degrees 96 * (pi/180) --> gives you radians

OpenStudy (anonymous):

so you have to divide by pi? so then it would be \[\frac{ 96 }{ \pi }\times 180?\]

OpenStudy (dumbcow):

?? why divide --> 96*(pi/180) = 96pi/180 reduce fraction

OpenStudy (anonymous):

so what would that be

OpenStudy (dumbcow):

cmon you can figure it out example: 8/12 = 4/6 = 2/3

OpenStudy (anonymous):

my calculator died i dont have any batteries...i really just want the answer